OpenStudy (anonymous):

I have seen around that the derivative of exponential functions like a^(x) is the function it self a^(x) times the derivative of the exponent (in this case is just 1) times ln(a). Is this correct for any kind of exponential function?

OpenStudy (phi):

the idea is you can write a as \[a= e^{some\_number} \] by definition, the some number is ln(a) that is \[ a= e^{\ln a} \] thus \[ a6x = \left( e^{\ln a}\right)^x = e^{x \ln a}\]

OpenStudy (phi):

*\(a^x\)

OpenStudy (phi):

ln a is a constant (call it c), and the derivative of exp( c x) is exp( c x) * d (c x) or c exp(c x)

OpenStudy (anonymous):

So, at the end we apply the definition of the derivative of "exp()", which actually I am not aware of, but apparently the derivative of "exp(x)" is exp(x) * (x)', right? But I am not seeing the connection between the function exp and e, are they the same?

OpenStudy (phi):

sorry, yes exp(x) means e^x (just easier to type)